Demonstrate a sophisticated understanding of effects of different compounding periods and interest rates on future value of money,
Answer #1

Compounding is a technique of time value of money, which is used for finding out the future value of investment. Future value is calculated with the help of interest rate offered on investment and horizon period for investment. As the interest rate will increase keeping the investment horizon period constant, future value of investment will increase and vice versa. For example You have invested 10000 in a bank at a interest rate of 5% for 5 year and now another bank offer same deposit at a rate of 6% then future value of both deposits would be different due to different interest rate.

Future value of deposit with 5% interest rate = PV*(1+r)^n = 10000*(1.05)^5 = 12762.82

Future value of deposit with 6% interest rate = PV*(1+r)^n = 10000*(1.06)^5 = 13382.26

Now take the above example in which period for investment has changed to 5 years for option 1 and 6 years for deposit two and rate of interest is same for both the deposits at 5%

Future value of deposit with 5% interest rate for 5 years = PV*(1+r)^n = 10000*(1.05)^5 =12762.82

Future value of deposit with 5% interest rate for 5 years = PV*(1+r)^n =10000*(1.06)^5 =13400.96

Thus from the above example we can say that any change in interest rate and horizon investment period will affect the future value of Investment.